after driving 100km 2/7 of the trip is completed how long is the total trip
Answer by jim_thompson5910(35256)

Let x = total distance

If 100 km represents 2/7 of the trip, then this means that 100 = (2/7)x, solve for x to get


100 = (2/7)x

100*7 = 2x

700 = 2x

700/2 = x

350 = x

x = 350


So the total distance is 350 km